OpenStudy (anonymous):

Find the exact value of cos^2(1/2cos^-1 5/13)

OpenStudy (aum):

Let \(\large \cos^{-1}\frac{5}{13} = x\). Then \(\large \cos(x) = \frac{5}{13}\) \[ \cos^2(1/2\cos^{-1} 5/13) = \cos^2(x/2) \\ \cos(2A) = 2\cos^2A - 1 \\ 2\cos^2(A) = \cos(2A) + 1 \\ \cos^2(A) = \frac 12 * (\cos(2A) + 1)\\ \cos^2(x/2) = \frac 12 * (\cos(x) + 1) = \frac 12 * \left ( \frac {5}{13}+1\right ) = \frac 12 * \frac {18}{13} = \frac{9}{13} \]